# Scalable Multi-view Clustering via Explicit Kernel Features Maps

### v1

@@ 4.4 Complexity Analysis
ACM DBLP Amazon Photos Amazon Computers OGBN-ArXiv OGBN-Products
LMVSC 2.36 ±0.1 2.97 ±0.1 37.78 ±3.7 144.26 ±14.2 2952.97 ±210.43 Timeout
@@
MvAGC 5.05 ±0.1 5.65 ±0.0 36.59 ±1.0 36.76 ±0.7 Out of Memory Out of Memory
MAGC 21.83 ±1.6 38.12 ±3.8 304.23 ±22.6 8699.82 ±16343.2 Out of Memory Out of Memory
@@
MvSCK 0.18 ±0.1 0.19 ±0.1 0.86 ±0.1 5.96 ±0.2 20.43 ±3.8 239.32 ±14.2
@@
ACM DBLP Photos Computers ArXiv Products
@@
Quadratic 93.2 0.2 93.1 0.2 75.7 0.9 66.3 6.0 48.3 20.4 47.1 239.3
RBF 93.2 0.8 93.1 2.0 76.2 2.4 65.0 6.5 48.7 15.8 18.6 325.2
Sigmoid 93.2 0.9 93.0 1.9 74.7 2.3 65.6 6.4 34.0 16.6 33.2 231.3
(a) ACM
(b) DBLP
(c) Amazon Photos
(d) Amazon Computers
(e) OGBN-ArXiv
(f) OGBN-Products
Figure 2: Sensitivity of MvSCK w.r.t the temperature parameter $T$. Note that in our experiments, we used $T=0.1$ for all datasets.
5 Experiments
  • How does MvSCK compare with state-of-the-art in terms of:
    • RQ1: Clustering performance?
    • RQ2: Computational efficiency?
  • RQ3: Does our approach consistently outperform the state-of-the-art ?
  • RQ4: How do the hyper-paramters affect MvSCK?
  • RQ5: Is the view-importance parameter necessary?
5.1 Datasets
We evaluate our method on six benchmark datasets with different degrees of overlap: two small scale datasets, namely ACM and DBLP [4]; two medium datasets, namely Amazon Photos and Amazon Computers [19]; and two large datasets OGBN-ArXiv and OGBN-Products [9]. The six datasets are described briefly below:
  • ACM is a network dataset of research papers, where the associated node features are the papers’ keywords in bag-of-words format. The “Co-Subject” and “Co-Author” graphs respectively connect papers with the same theme or the same author.
  • DBLP is a network dataset of authors with bag-of-words node features representing keywords. “Co-Author”, “Co-Conference”, and “Co-Term” links indicate that the authors concerned have collaborated on papers, published papers at the same conference, or published papers with the same terms.
  • Amazon Photos & Amazon Computers are segments of the Amazon co-purchase network dataset. The nodes represent goods, and their attribute features are represented by bag-of-words of product reviews. The edges indicate that two goods are purchased together. The second set of attributes represents the word co-occurrence matrix.
  • OGBN-ArXiv represents a citation network encompassing a set of computer science arXiv papers. Each paper comes with a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract. We add a second $k$-nearest neighbors (k-NN) graph with 10 neighbors and added self-connections.
  • OGBN-Products is a graph representing an Amazon product co-purchase network. An edge between two products signifies that those products are purchased together. The node features are created by extracting bag-of-words features from the product descriptions and reducing their dimensionality to 100 via PCA. Similarly, to OGBN-ArXiv, we add a second k-NN graph.
5.2 Experimental Setup
Experimental setting
For the purposes of comparison, we validate our approach against the state-of-the-art approaches discussed in the Related Work section above, with reference to clustering accuracy (CA), clustering F1-score (CF1), adjusted rand index (ARI), and normalized mutual information (NMI). We use the official implementations of compared approaches. Parameters are set based on the recommendations of the authors. When these recommendations are not available, we set parameters similarly to how they were set on datasets of similar dimensions. To make comparisons fair we remove any supervised signals when running experiments in the original code; these would include, for example, the validation of runs based on supervised metrics such as accuracy or setting random seeds. For our approach, the kernel used is the quadratic kernel that was discussed in [5]. The number of components parameter $f$ was set equal to the number of clusters on all datasets, with the exception of OGBN-Products, where it was set to 10 due to its large size. Note that all metrics reported are averaged over five different runs. All experiments were performed on a laptop equipped with a 12th Gen Intel(R) Core(TM) i9-12950HX (24 CPU threads) 2.3 GHz processor, a 12GB GeForce RTX 3080 Ti GPU, and 64GB of RAM.
Learning Node Representations
To perform clustering on network datasets using generic approaches such as ours in addition to LMVSC, SMVSC, FDAGF and OMSC, we first need to learn node embeddings. Given a multi-view attributed network consisting of adjacency matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{V}$ and feature matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, we learn embeddings via Laplacian smoothing. This is a well-known technique that has been used in works such as [17, 6] to handle graph data. It involves consecutive neighborhood averaging steps:
where $p$ is the propagation order. Specifically, in our experiments, we set the propagation order as follows: 2 for ACM, 2 for DBLP, 20 for Amazon Photos, 40 for Amazon Computers, 60 for OGBN-ArXiv, and 100 for OGBN-Products. This step can be seen as a preprocessing step that is independent of the approaches themselves. It is a methodology that has emerged from work done on simplifying graph convolutional networks [27].
5.3 Clustering Results (RQ1)
The clustering outcomes for small, medium, and large graph datasets are detailed in Tables 3, 4, and 5 respectively. A comprehensive analysis of these tables reveals the superior performance of our approach relative to other methodologies across a majority of datasets and evaluation metrics. Remarkably, our method exhibits superiority even when compared to specialized techniques designed explicitly for graph datasets. Noteworthy observations include LMVSC’s competitiveness with our approach on the Amazon Computers dataset, where both methods achieve comparable results. Basically, while LMVSC outperforms our approach in terms of F1 and ARI metrics, our method outperforms it in accuracy and NMI metrics. This comparison underscores the strengths of our approach, showcasing its robust performance across diverse datasets and metrics. We observed some disparities between results reported in original papers and those obtained by running the code on the same datasets. Additionally, many approaches struggle to scale to challenging datasets like OGBN-ArXiv and OGBN-Products. Specifically, OMSC fails to retrieve the correct number of clusters in Amazon Photos and Computers, impacting its ability to compute CF1 scores.
5.4 Running Times (RQ2)
The detailed breakdown of running times for different models across all datasets is presented in Table 6, where our approach emerges as the frontrunner with markedly shorter execution times across small, medium, and large graph datasets. The notable efficiency of our method becomes especially evident when compared to alternative approaches. Notably, for the challenging OGBN-ArXiv and OGBN-Products datasets, our approach not only maintains its speed advantage but also adheres to the strict three-hour time limit we set. In contrast, several competing models exceed this limit, highlighting the computational efficiency of our proposed method in the face of diverse datasets.
5.5 Statistical Significance Testing (RQ3)
Figure 4 shows the results of the Holm mean-rank post-hoc test [8] over the clustering results. We compute the mean rank over rankings generated based on each (dataset, metric, run) triplet. We see that our approach is ranked first by a large margin, followed by LMGEC and LMVSC as second and third. After these three come various overlapping groups of approaches.
In Figure 5, we report the results of the Holm mean-rank post-hoc test over the execution times of the different algorithms. Ours is the most efficient, followed by LMGEC. LMVSC and MvAGC are ranked third, followed by SMVSC, OMSC, and then a group composed of MAGC and FDAGF. MCGC comes last.
Figure 3: Clustering accuracy with and without the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$.
5.6 Sensitivity Analysis (RQ4)
Temperature Parameter $\ T$
In our experiments we tried various values for the temperature hyperparameter. Figure 2 shows the performance of LMGEC for different temperature values on the different datasets and for the different clustering metrics. We see that on most datasets the performance remains mostly constant but that larger values deteriorate the performance. As such, we advise using $T\leq 1$. In our exeperiments, we set it to $T=0.1$ for all datasets.
Kernel Feature Map
In table 7 we report the results of our approach using different kernels, namely an exact quadratic kernel, Nystroem approximations of the radial basis function (RBF), and a sigmoid kernel. For the approximations we use $10*k$ components, apart from on the OGBN-Products dataset, which is too large and for which we use $100$ components only. The results show that using different kernels gives similar performances, except on OGBN-Products, where the dimension of the approximations would appear to be insufficient, leading to a poor approximation of the feature maps; the same is true for the sigmoid kernel on ArXiv. Note that in the context of RBF and sigmoid we use approximations, given that their exact feature maps are infinite-dimensional.
@@ 5.7 Ablation relative to the View-Importance Parameter $\boldsymbol{\lambda}$ (RQ5)
Figure 3 illustrates the impact of the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$ on various datasets. The influence of $\boldsymbol{\lambda}$ is evident in its ability to enhance performance across five out of the six datasets analyzed during ablation studies. Notably, there is no observed deterioration in performance on the remaining ACM dataset, indicating that the regularization vector does not adversely affect outcomes in this particular case.
11113333555577779999MvSCKLMGECLMVSCSMVSCFDAGFMAGCOMSCMCGCMvAGC
11113333555577779999MvSCKLMGECLMVSCMvAGCSMVSCOMSCMAGCFDAGFMCGC

### v2

@@ 6.4 Running Times (RQ2)
ACM
DBLP
Photos
Computers
ArXiv
Products
LMVSC 2.36 ±0.1 2.97 ±0.1 37.78 ±3.7 144.26 ±14.2 2952.9 ±210.4 Timeout
@@
MvAGC 5.05 ±0.1 5.65 ±0.0 36.59 ±1.0 36.76 ±0.7 OOM OOM
MAGC 21.83 ±1.6 38.12 ±3.8 304.23 ±22.6 8699.82 ±16343.2 OOM OOM
@@
MvSCK
0.18 ±0.1
0.19 ±0.1
0.86 ±0.1
5.96 ±0.2
20.43 ±3.8
239.32 ±14.2
6.5 Statistical Significance Testing (RQ3)
Figure 2 shows the results of the Holm mean-rank post-hoc test [holm1979simple] over the clustering results. We compute the mean rank over rankings generated based on each (dataset, metric, run) triplet. We see that our approach is ranked first by a large margin, followed by LMGEC and LMVSC as second and third. After these three, there are various overlapping groups of approaches.
In Figure 3, we report the results of the Holm mean-rank post hoc test on the execution times of the different algorithms. Ours is the most efficient, followed by LMGEC. LMVSC and MvAGC are ranked third, followed by SMVSC, OMSC, and then a group composed of MAGC and FDAGF. MCGC comes last.
1133557799
MvSCK
LMGECLMVSCSMVSCFDAGFMAGCOMSCMCGCMvAGC
1133557799
MvSCK
LMGECLMVSCMvAGCSMVSCOMSCMAGCFDAGFMCGC
6.6 Sensitivity Analysis (RQ4)
ACM
DBLP
Photos
Computers
ArXiV
Products
CA
CF1
NMI
ARI
Figure 5: Sensitivity of MvSCK in terms of CA, CF1, NMI and ARI according to the temperature parameter $T$.
6.6.1 Temperature Parameter T\ T
In our experiments we tried various values for the temperature hyperparameter. Figure 5 shows the performance of MvSCK for different temperature values in different data sets and for the different clustering metrics. We see that on most datasets the performance remains mostly constant, but that larger values deteriorate the performance. As such, we advise using $T\leq 1$. In our experiments, we set it to $T=0.1$ for all datasets.
6.6.2 Kernel Feature Map
In Table 7 we report the results of our approach using different kernels, namely an exact quadratic kernel, Nystroem approximations of the radial basis function (RBF), and a sigmoid kernel. For the approximations we use $10\times k$ components, apart from on the Products dataset, which is too large and for which we use $100$ components only. The results show that using different kernels gives similar performances, except on Products, where the dimension of the approximations would appear to be insufficient, leading to a poor approximation of the feature maps; the same is true for the sigmoid kernel on ArXiv. Note that in the context of RBF and sigmoid we use approximations, given that their exact feature maps are infinite-dimensional.
@@ 6.6.2 Kernel Feature Map
ACM
DBLP
Photos
Computers
ArXiv
Products
@@
Quadratic
93.2
0.2
93.1
0.2
75.7
0.9
66.3
6.0
48.3 20.4
47.1
239.3
RBF
93.2
0.8
93.1
2.0
76.2
2.4 65.0 6.5
48.7
15.8
18.6 325.2
Sigmoid
93.2
0.9 93.0 1.9 74.7 2.3 65.6 6.4 34.0 16.6 33.2
231.3
@@ 6.7 Ablation relative to the View-Importance Parameter $\boldsymbol{\lambda}$ (RQ5)
Figure 6 shows the effect of the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$ on the different datasets. $\boldsymbol{\lambda}$ is seen to boost performance on five out of the six datasets that we considered for ablation, and does not impair performance on the remaining dataset ACM. Figure 6 illustrates the impact of the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$ on various datasets. The regularization vector $\boldsymbol{\lambda}$ improves performance in five of the six datasets analyzed during ablation studies. Notably, there is no deterioration in performance on the remaining ACM dataset, indicating that the regularization vector does not adversely affect the outcome in this particular case.
Figure 6: Clustering accuracy with and without the regularization vector $\boldsymbol{\lambda}$